OpenStudy (anonymous):

How would you go about finding the solution to this system of three equations in 3 variables?

OpenStudy (anonymous):

can i see those three equations it might help

OpenStudy (anonymous):

2x - y + z = 7 x + 2y - 5z = -1 x - y = 6

OpenStudy (anonymous):

you can probably use the last equation and make x = y + 6 substitute that value into the first or second equation and find y = and then substitue that back into the last equation to get x = i did this 3 years ago in my freshman year but you can try it. If you want i will try to figure this question out and tel you the answer 10 minutes from now? :O

OpenStudy (anonymous):

its okay, you don't have to work it out hun (:

OpenStudy (anonymous):

just to be on the safe side :)

OpenStudy (anonymous):

Here it goes: Label all the equations by numbers...gonna be useful 2x - y + x =7 (1) x + 2y - 5z = -1 (2) x - y = 6 (3) find a equation with a coefficient you can eliminate using another equation I chose (1) and (2) Make one variable have the same coefficient Since the third equation consist only of x and y --> we need to eliminate z in this one 2x - y + z = 7 (1) x 5 x + 2y -5z = -1 (2) 10x - 5y + 5z = 35 + x + 2y -5z = -1 ------------------- 11x - 3y = 34 Now that you made an equation consisting of x and y: solve for one variable using substitution or elimination rearrange (3) x - y = 6 x = y + 6 substitute x = y - 6 into 11x - 3y = 34 11x - 3y = 34 11(y + 6) - 3y = 34 11y + 66 - 3y = 34 * combine the like terms and move 66 to the right side 11y - 3y = 34 - 66 8y = -32 * Divide both sides by 8 y = -4 one variable down two more to go substitute y = -4 into (3). Make sure you use the original equation and not the re arranged one :D x - y = 6 x - (-4) = 6 x + 4 = 6 * Move 4 over to the other side x = 6 - 4 x = 2 Alright now you got two variables and you need the last one :D substitute x = 2 and y = -4 into equation (2) 2x - y + z = 7 (2) 2(2) - (-4) + z = 7 4 + 4 + z = 7 8 + z = 7 z = 7 - 8 z = -1 You got all three variables. Write your therefore statement. Therefore, x = 2, y = 4 and z = -1 Moral of the question: the order matters

OpenStudy (anonymous):

thank you (: